Jill is 20 km away from joe. Jill walks at 2.5km/hr they meet in 5 hrs how fast is joe walking
Answer by stanbon(75887)

Jill is 20 km away from joe. Jill walks at 2.5km/hr
they meet in 5 hrs how fast is joe walking.
-------------
Jill DATA:
time = 5 hrs ; rate = 2.5 km/hr ; distance = 5*2.5 = 12.5 km
--------------
Joe DATA:
time = 5 hrs ; distance = 20-12.5 = 7.5 km ; rate = 7.5/5 = 1.5 mph
--------------------
Cheers,
Stan H.